Consider an economy described by the following equations:

 Y=C+I+GY=C+I+G C=100+0.75×(Y - T)C=100+0.75×Y - T I=500?50×rI=500?50×r G=125G=125 T=100T=100

where Y is GDP, C is consumption, II is investment, G is government purchases, T is taxes, and r is the interest rate. If the economy were at full employment (that is, at the natural rate of output), GDP would be $2,000.

Identify the equation(s) each of the following statements describes. Check all that apply.

StatementCIGTIt is an autonomous amount, independent of other factors.     It is a function of disposable income.     It depends on the interest rate.     

The marginal propensity to consume in this economy is ___

Suppose the central bank’s policy is to adjust the money supply to maintain the interest rate at 4%, so r=4.

When the interest rate is 4%, GDP is ___

True or False: GDP at an interest rate of 4% is below the full-employment level.

Assuming no change in monetary policy, (a decrease, an increase) in government purchases by ____ would restore GDP to the full-employment level. (Note: Assume that such change in fiscal policy has no crowding-out effect.)

Assuming no change in fiscal policy, (a decrease. an increase) in the interest rate by ___ would restore GDP to the full-employment level.
Answer #1

It is an autonomous amount, independent of other factors - G and T

It is a function of disposable income - C

It depends on the interest rate - I

The marginal propensity to consume in this economy is = 0.75 (coefficient of Yd = Y - T in the equation for C)

r = 4

I = 300

C + I + G = Y

525 + 0.75 x (Y - 100) = Y

GDP = Y = 1800

The statement is true (GDP < potential GDP = 2000)

output gap = 2000 - 1800 = 200

increase in GDP x 1/(1 - MPC) = 200

Assuming no change in monetary policy, an increase in government purchases by 50 would restore GDP to the full-employment level.

Assuming no change in fiscal policy, a decrease in the interest rate by 1% would restore GDP to the full-employment level. (r = 3 would mean I = 350 and Y = 2000)